A rectangular block of ice with dimensions 2 m by 2 m by 0.3 m floats on water. A person weighing 835 N wants to stand on the ice. Would the ice sink below the surface of the water?



1
Expert's answer
2021-11-23T10:54:40-0500

If "F_B" (Buoyant Force) "\ge" (m+M) g

Then the ice will float.

"F_B=(\rho_wV_{ice}) g"

"=(1000\u00d72\u00d72\u00d70. 2)9.8"

"=7840N"


"mg+Mg=980+(\rho_{ice}V_{ice} g)"

"=(980+(916.7\u00d72\u00d72 \u00d70. 2\u00d79.8)"

"=980+7186.93"

"=8166.93N \gt F_B"


The ice will sink since "F_B" is not greater than "(m+M) g"